### v2

# Balancing Embedding Spectrum for Recommendation

Abstract.
Modern recommender systems heavily rely on high-quality representations learned from high-dimensional sparse data. While significant efforts have been invested in designing powerful algorithms for extracting user preferences, the factors contributing to good representations have remained relatively unexplored. In this work, we shed light on an issue in the existing pair-wise learning paradigm (i.e., embedding collapse), that the representations tend to span a subspace of the whole embedding space, leading to a suboptimal solution and reducing the model capacity. Specifically, we show that alignment of positive pairs is equivalent to a low pass filter causing users/items to collapse to a constant vector. While negative sampling can partially mitigate this issue by acting as a high pass filter to balance the spectrum, leading to an incomplete collapse.
To tackle this issue, we present a novel learning paradigm DirectSpec, which directly optimizes the spectrum distribution to ensure that users/items effectively span the entire embedding space. We demonstrate that many self-supervised learning algorithms without explicit negative sampling can be considered as special cases of DirectSpec. Furthermore, we show that optimizing the spectrum inappropriately could also be poisonous to data representation, where the key lies in a dynamic balance between alignment of positive pairs and spectrum balancing. Finally, we propose an enhanced and practical implementation DirectSpec+ to balance the embedding spectrum more adaptively and effectively. We implement DirectSpec+ on two popular recommender models: MF and LightGCN. Our experimental results demonstrate its effectiveness and efficiency over competitive baselines.
Recommender System, Collaborative Filtering, Embedding Collapse, Embedding Spectrum
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Information systems Recommender systems
1. Introduction
Recommender systems have penetrated into our daily life, we can see them everywhere such as e-commerce (Wang et al., 2018), short-video (Liu et al., 2019), social network (Jiang et al., 2012), and so on. Collaborative filtering (CF), a fundamental technique for recommendation to discover user preference based on the historical data, has attracted much attention in the last decades. The most extensively used CF technique, matrix factorization (MF) (Koren et al., 2009) which represents users and items as low dimensional latent vectors and estimates the rating as the inner product between latent vectors, has been the cornerstone of modern recommender systems. Since MF estimates the rating with a simple linear function, subsequent works mostly focus on designing more powerful and complex algorithms to model non-linear user-item relations, including but not limited to multi-layer perceptron (MLP) (He et al., 2017), attention mechanism (Kang and McAuley, 2018), reinforcement learning (Zheng et al., 2018), transformer (Sun et al., 2019), diffusion model (Wang et al., 2023), graph neural network (GNN) (Ying et al., 2018; Peng et al., 2022b), etc., and have shown tremendous success.
Although different kinds of methods have been proposed, most of them can be considered as MF variants whose goal is to learn low dimensional representations (with dimension $d$) from the high dimensional sparse interaction matrix (with dimension $D\gg d$). Figure 1 illustrates the top 500 normalized singular value distribution of the interaction matrix of CiteULike (see Section 5.1 for data description). We observe that users/items are predominantly distributed along only a few dimensions in the original space while most dimensions barely contribute (i.e., with singular values close to 0) to the representations. Thus, when users and items are mapped into a more compact embedding space, it is expected that redundant dimensions are all removed and each dimension contributes to the user/item representations as equally and uniformly as possible (i.e., the representations make full use of the embedding space).
Existing recommendation methods mostly utilize positive (observed) and negative (unobserved) samples for optimization. Here, a simple yet fundamental question arises: Can user/item representations of existing works make full use of all dimensions? Unfortunately, by analyzing the spectrum of the embedding matrix, we empirically and theoretically show that users/items tend to span a subspace of the whole embedding space (with dimension $d^{\prime}<d$), where the embeddings collapse along all (complete collapse) or certain dimensions (incomplete collapse). Particularly, alignment of positive pairs is equivalent to a low pass filter, where the representations tend to collapse to a constant vector. Negative sampling is commonly used to optimize recommendation algorithms without causing an explicit embedding collapse by pushing away the unobserved user-item pairs, and we show that it is equivalent to a high pass filter decelerating the speed of collapse by balancing the embedding spectrum. However, the effectiveness of negative sampling is only limited to certain loss functions such as Bayesian personalized ranking (BPR) (Rendle et al., 2009) and binary cross-entropy (BCE) loss (He et al., 2017), where collapse over certain dimensions still exists.
Figure 1. An illustration of the data distribution in the original space and expected distribution in the learning embedding space.
In this work, we tackle the embedding collapse issue from a spectral perspective. We observe that the extent of the collapse is closely related to the spectrum distribution of the embedding matrix. Specifically, only a single singular value dominates when the representation completely collapses, whereas the singular values are uniformly distributed when the representations make full use of the embedding space. Inspired by this observation, we propose to directly optimize the spectrum distribution (dubbed DirectSpec) ensuring that all dimensions equally contribute to the representations. We prove that DirectSpec can completely prevent the embedding collapse and provide a practical implementation with a complexity as $\mathcal{O}(B^{2}d)$ where $B$ is the batch size. In addition, we demonstrate that advanced self-supervised learning (SSL) algorithms such as Barlow Twins, LogDet, and spectral contrastive loss without explicit negative sampling can be considered as special cases of DirectSpec. Furthermore, we show that optimizing the spectrum distribution could also be poisonous to data representation, where the key lies in a dynamic balance between alignment of positive pairs and spectrum balance. Based on this observation, we point out the limitations of DirectSpec and propose an enhanced variant DirectSpec+ to balance the embedding spectrum more adaptively and effectively. Finally, we implement DirectSpec and DirectSpec+ on two popular baselines: MF (Koren et al., 2009) and LightGCN (He et al., 2020), and experimental results show that DirectSpec+ improves BPR and LightGCN by up to 52.6% and 42.4% in terms of nDCG@10, respectively. The contributions of this work can be summarized as follows:
  • We theoretically and empirically show that existing recommendation methods suffer from embedding collapse, where the representations tend to fall into a subspace of the whole embedding space, and analyze the mechanisms causing this issue.
  • We propose a novel learning paradigm DirectSpec which directly optimizes the spectrum distribution, ensuring that users/items effectively span the entire embedding space. We empirically and theoretically show that DirectSpec can prevent embedding collapse.
  • We demonstrate that SSL methods such as Barlow Twins, LogDet, and spectral contrastive loss without explicit negative sampling can be considered as special cases of DirectSpec.
  • Extensive experimental results on three datasets demonstrate the efficiency and effectiveness of our proposed methods.
2. Preliminaries
2.1. Pairwise Learning for CF
Given the user set $\mathcal{U}$ and item set $\mathcal{I}$, the interaction matrix is defined as $\mathbf{R}\in\{0,1\}^{\left|\mathcal{U}\right|\times\left|\mathcal{I}\right|}$, the observed interactions are represented as $\mathbf{R}^{+}=\{r_{ui}=1|u\in\mathcal{U},i\in\mathcal{I}\}$. Users and items are first mapped to low dimensional vectors $\mathbf{H}\in\mathbb{R}^{(\left|\mathcal{U}\right|+\left|\mathcal{I}\right|)\times d}$ through an encoder, where the encoder can be as simple as a linear mapping (Koren et al., 2009) or advanced algorithms such as MLPs (He et al., 2017), GNNs (Wang et al., 2019), etc. Let $\mathbf{h}_{u}$ and $\mathbf{h}_{i}$ be the $u$’s and $i$’s representations, respectively. The goal of CF is to predict unobserved interactions $\mathbf{R}^{-}=\{r_{ui}=0|u\in\mathcal{U},i\in\mathcal{I}\}$ estimated as the inner product between the user and item representations: $\hat{r}_{ui}=\mathbf{h}_{u}^{T}\mathbf{h}_{i}$. The model parameters $\Theta$ are optimized through a loss function $\mathcal{L}$ formulated as follows:
(1) ${\rm arg}\mathop{{\rm min}}\limits_{\mathbf{\Theta}}\mathcal{L}(\hat{r}_{ui},r_{ui}).$
The loss function measures the difference between the estimated score and the ground truth. For instance, BPR and BCE loss are extensively used for CF methods:
(2) $\mathcal{L}_{BPR}=\sum_{(u,i)\in{\mathbf{R}}^{+},(u,j)\in{\mathbf{R}}^{-}}-\ln\sigma(\hat{r}_{ui}-\hat{r}_{uj}),$
$\mathcal{L}_{BCE}=\sum_{u\in\mathcal{U},i\in\mathcal{I}}-r_{ui}\ln\sigma(\hat{r}_{ui})-\left(1-r_{ui}\right)\ln\left(1-\sigma(\hat{r}_{ui})\right),$
[…]
Definition 0.
Effective Rank. Given singular values of the embedding matrix $\mathbf{H}$: $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{d}\geq 0$, let $p_{k}=\frac{\sigma_{k}}{\sum_{k}\sigma_{k}}$, then the effective rank is defined as follows:
(3) $erank(\mathbf{H})=exp\left(H\left(p_{1},\cdots,p_{d}\right)\right),$
where $H\left(p_{1},\cdots,p_{d}\right)=-\sum_{k}p_{k}\log p_{k}$ is the Shannon entropy.
Compared with rank, erank takes the singular value distribution into consideration: the more uniform (sharper) of the distribution, the higher (lower) of the erank (Roy and Vetterli, 2007). The embedding matrix contains the least information when there is only one leading non-zero singular values ($\sigma_{2}=\cdots=\sigma_{d}=0$, and $erank(\mathbf{H})=rank(\mathbf{H})=1$), while erank is maximized when each dimension equally contributes to the representations: $\sigma_{1}=\cdots=\sigma_{d}$ ($erank(\mathbf{H})=rank(\mathbf{H})=d$).
2.2. Graph Representation and Graph Filtering
The interactions can be represented as a graph. Consider a graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{A})$, where the node set contains all users and items: $\mathcal{V}=\mathcal{U}\cup\mathcal{I}$, the edge set is represented by observed interactions: $\mathcal{E}=\mathbf{R}^{+}$, $\mathbf{A}$ is an adjacency matrix. Suppose $\mathbf{A}$ is normalized with the eigenvalue $|\lambda_{k}|\leq 1$, then:
Definition 0.
Graph Filtering. Let $11=\lambda_{1}>\lambda_{2}>\cdots>\lambda_{n}=-$ 1 be the eigenvalues of $\mathbf{A}$ where $n=|\mathcal{V}|$. The components closer to $\lambda_{1}$ and $\lambda_{n}$ correspond to a lower and higher frequency, respectively. $\mathcal{F}^{L}(\mathbf{A})$ is a low pass filter if $|\mathcal{F}^{L}(\lambda_{i})|>|\mathcal{F}^{L}(\lambda_{j})|$ for $\lambda_{i}>\lambda_{j}$ and $\mathcal{F}^{H}(\mathbf{A})$ is a high pass filter if $|\mathcal{F}^{H}(\lambda_{i})|>|\mathcal{F}^{H}(\lambda_{j})|$ for $\lambda_{i}<\lambda_{j}$.
Note that $\mathcal{F}^{L}(\mathbf{A})$ is defined as a low pass filter as it emphasizes more on the lower frequencies and similarly for the definition of the high pass filter. Please refer to (Sandryhaila and Moura, 2014) for detailed introduction of graph frequency.
Figure 2. (a) complete collapse, (b) incomplete collapse, (c) how negative sampling can balance the spectrum on Yelp: the normalized eigenvalue distribution (Top 500) of $\mathbf{A}-\mathbf{\bar{A}}$.
3. Embedding Collapse in CF
3.1. Complete Collapse
Let us first consider a generic loss function only optimizing alignment of positive pairs (i.e., minimizing their distance): $\mathcal{L}(\hat{r}_{ui},r_{ui}),(u,i)\in\mathbf{R}^{+}$.
Proposition 1.
Suppose $G$ is connected, then $\mathbf{h}_{k}\approx\mathbf{h}_{z}$ for arbitrary nodes $k$ and $z$ when $\mathcal{L}$ completely converges.
By calculating the gradient over the embeddings, the parameters are updated through stochastic gradient descent (SGD) as follows:
(4) $\mathbf{H}^{(l+1)}=\mathbf{H}^{(l)}-\alpha\frac{\partial\mathcal{L}}{\partial\mathbf{H}^{(l)}}=\mathbf{H}^{(l)}+\alpha\mathbf{A}\mathbf{H}^{(l)},$
$\mathbf{h}^{(l+1)}_{u}=\mathbf{h}^{(l)}_{u}-\alpha\frac{\partial\mathcal{L}}{\partial\mathbf{h}_{u}^{(l)}}=\mathbf{h}^{(l)}_{u}+\alpha\sum_{(u,i)\in\mathbf{R}^{+}}\mathbf{A}_{ui}\mathbf{h}^{(l)}_{i},$
where $\alpha\in(0,1)$, $\mathbf{A}_{ui}=-\frac{\partial\mathcal{L}}{\partial\hat{r}_{ui}}$ for $(u,i)\in\mathbf{R}^{+}$. Ideally, if $\frac{\partial\mathcal{L}}{\partial\hat{r}_{ui}}$ remain unchanged during training such as the Euclidean distance loss: $\mathcal{L}=\sum\left\|\mathbf{h}_{u}-\mathbf{h}_{i}\right\|^{2}_{2}$, then Equation (4) can be rewritten as:
(5) $\mathbf{H}^{(l)}=\left(\mathbf{I}+\alpha\mathbf{A}\right)\mathbf{H}^{(l-1)}=\cdots=\left(\mathbf{I}+\alpha\mathbf{A}\right)^{l}\mathbf{H}^{(0)}.$
Suppose $\mathbf{A}$ is normalized such as $|\lambda_{k}|\leq 1$, then $(\mathbf{I}+\alpha\mathbf{A})^{l}$ is equivalent to a low pass filter, as the eigenvalue of $\mathbf{I}+\alpha\mathbf{A}$ is $1+\alpha\lambda_{k}\in[1-\alpha,1+\alpha]$ with $01-\alpha>$ 0 and $x^{l}$ is monotonically increasing for $l>0$ and $x>0$. According to the spectral decomposition with $\mathbf{v}_{k}$ as the eigenvector:
(6) $\left(\mathbf{I}+\alpha\mathbf{A}\right)^{l}=\sum_{k}(1+\alpha\lambda_{k})^{l}\mathbf{v_{k}}\mathbf{v_{k}}^{T},$
it is obvious that $\frac{(1+\alpha\lambda_{k})^{l}}{(1+\alpha\lambda_{1})^{l}}\rightarrow 0$ ($k\neq 1$) as $l$ is large enough, resulting in rank$((\mathbf{I}+\alpha\mathbf{A})^{l})\rightarrow 1$, then:
(7) ${\rm rank}\left(\mathbf{H}^{(l)}\right)\leq{\rm min}\left({\rm rank}\left((\mathbf{I}+\alpha\mathbf{A})^{l}\right),{\rm rank}\left(\mathbf{H}^{(0)}\right)\right)=1,$
[…]
  • Before the training starts, the embeddings are uniformly distributed in the embedding space since $erank(\mathbf{H})\approx d$.
  • The erank monotonically declines on both models, indicating that the embeddings collapse as the training starts and the issue becomes more serious as training proceeds.
  • The erank decreases more rapidly on LightGCN than MF. Consider a $K$-layer LightGCN, then Equation (5) can be rewritten as:
    (8) $\mathbf{H}^{(l)}=\left(\mathbf{I}+\alpha\mathbf{A}\right)^{l}\sum_{k=0}^{K}\mathbf{A}^{k}\mathbf{E}.$
    Here, $\mathbf{E}$ is the initial stacked user/item embeddings sent to the encoder; we ignore the difference between the adjacency matrix used in Equation (5) and LightGCN for simplicity. We can see that increasing the layer $K$ causes the loss function to converge faster, indicating that the over-smoothing in GNN (Li et al., 2018) aggravates the collapse issue.
  • The erank tends to converge to a value larger than 1 due to the two reasons we mentioned above. By comparing the Euclidean distance loss (MF) and log loss (MF), we can see the erank of the Euclidean distance loss with stable gradients drops more quickly than that of log loss, which further verifies our analysis.
3.2. Incomplete Collapse
[…]
(9) $\mathbf{H}^{(l+1)}=\mathbf{H}^{(l)}+\alpha\left(\mathbf{A}-\mathbf{\bar{A}}\right)\mathbf{H}^{(l)},$
$\mathbf{h}^{(l+1)}_{u}=\mathbf{h}^{(l)}_{u}+\alpha\sum_{(u,i)\in\mathbf{R}^{+}}\mathbf{A}_{ui}\mathbf{h}^{(l)}_{i}-\alpha\sum_{(u,j)\in\mathbf{R}^{-}}\mathbf{\bar{A}}_{uj}\mathbf{h}^{(l)}_{j},$
where $\mathbf{\bar{A}}_{uj}=\frac{\partial\mathcal{L}}{\partial\hat{r}_{uj}}$ for $(u,j)\in\mathbf{R}^{-}$. Note that the gradients for positive and negative samples always have the opposite direction. Here, $\mathbf{I}+\alpha\left(\mathbf{A}-\mathbf{\bar{A}}\right)$ can be disentangled to a low pass $\mathbf{I}+\alpha\mathbf{A}$ and a high pass filter $\mathbf{I}-\alpha\mathbf{\bar{A}}$, where a high pass filter can balance the embeddings spectrum by reducing the weights on low frequency and raising the importance of high frequency components, leading to a more uniform distribution. We visualize the spectrum of $\mathbf{A}-\mathbf{\bar{A}}$ in Figure 2 (c), where we fix the elements of $\mathbf{A}$ and $\mathbf{\bar{A}}$ to 1 and normalize them separately. $\mathbf{\bar{A}}=0$ (zero matrix) for ‘Only positive’, and $\mathbf{\bar{A}}$ contains all unobserved interactions for ‘With negative’. It is obvious that $\mathbf{A}-\mathbf{\bar{A}}$ has a more balanced spectrum than $\mathbf{A}$, leading to a higher erank.
Proposition 2.
Negative sampling alone leads to collapse.
Suppose we only focus on the negative sampling with the updating rule as: $\mathbf{H}^{(l+1)}=\mathbf{H}^{(l)}-\alpha\mathbf{\bar{A}}\mathbf{H}^{(l)}$. Ideally (i.e., with the gradients remaining unchanged), the updating rule can be rewritten as follows:
(10) $\mathbf{H}^{(l)}=\left(\mathbf{I}-\alpha\mathbf{\bar{A}}\right)^{l}\mathbf{H}^{(0)}.$
Similar to the analysis in Section 3.1, the high frequency components gradually dominate ($\frac{(1-\alpha\lambda_{k})^{l}}{(1-\alpha\lambda_{n})^{l}}\rightarrow 0$) as $l$ becomes larger, and eventually ${\rm rank}\left(\mathbf{H}^{(l)}\right)\rightarrow 1$. Then, another question arises: If both positive and negative sampling lead to collapse, how do existing methods avoid the collapse?
We further zoom in on the gradient over BCE loss. The gradient over the positive sample $\mathbf{A}_{ui}$: $1-\sigma(\hat{r}_{ui})$ is close to 1 at the beginning of training, causing the embeddings to collapse. Since positive pairs are more likely to be sampled due to the data sparsity, the gradients vanish faster as they are pulled close: $\mathbf{A}\rightarrow 0$, allowing the high pass filter to balance the spectrum. As the negative samples are sampled and pushed away, we have: $\mathbf{\bar{A}}_{uj}=\sigma(\hat{r}_{uj})\rightarrow 0$, making the representations converge: $\mathbf{H}^{(l+1)}=\mathbf{H}^{(l)}$ and leading to an incomplete collapse. For BPR loss, since the magnitude of gradients is $1-\sigma(\hat{r}_{ui}-\hat{r}_{uj})$ for both positive and negative pairs, positive pairs being optimized causes the gradients for negative pairs to vanish as well, making the high pass filter have a weaker effect balancing the spectrum and also explaining why BPR(MF) converges more slowly than BCE(MF). To conclude, negative sampling is not designed to prevent the collapse issue, gradient vanishing on log loss is the ‘X factor’ to alleviate the complete collapse, and is also the reason causing the incomplete collapse as well. Therefore, a more effective design proved to completely solve the collapse issue is required.
4. Methodology
4.1. Directly Balancing Embedding Spectrum (DirectSpec)
The spectrum distribution can directly reflect the extent of embedding collapse. Intuitively, a uniform distribution indicates that different dimensions equally contribute to the embeddings and leads to a high effective rank. To this end, we consider the following Lagrangian:
(11) $\mathcal{L}=f\left(\sigma_{1},\cdots,\sigma_{d}\right)+\beta\left(\sqrt{\sum_{i}\sigma_{i}^{2}}-c\right),$
with the minimum reached at $\sigma_{1}=\cdots=\sigma_{d}=\frac{c}{\sqrt{d}}$, where the singular value is bounded by the $L_{2}$ ball with radius $c$.
Proposition 1.
Let $\frac{\partial f}{\partial\sigma_{i}}=g_{i}(\cdot)$. The minimum of Equation (11) is achieved at $\sigma_{1}=\cdots=\sigma_{d}=\frac{c}{\sqrt{d}}$ as long as $g_{i}(\sigma_{j})=g_{j}(\sigma_{j})$ for $i,j\in\{1,\cdots,d\}$.
By calculating the derivative over $\sigma_{i}$, we get:
(12) $\left\{\begin{array}[]{lr}g_{1}(\sigma_{1})+2\beta\sigma_{1}=0\\ \\ \qquad\;\;\cdots\\ \\ g_{d}(\sigma_{d})+2\beta\sigma_{d}=0\\ \\ s.t.\sum_{i}\sigma_{i}^{2}=c^{2}.\\ \end{array}\right.$
Obviously, $\sigma_{1}=\cdots=\sigma_{d}$ as long as $g_{i}(\cdot)$ has the same form for all singular values. For instance, a simple design is $f\left(\sigma_{1},\cdots,\sigma_{d}\right)=\sum_{i}\sigma_{i}$ where $\beta=-\frac{\sqrt{d}}{2c}$.
While the algorithm above has a provable guarantee to prevent the collapse issue, the efficiency is limited as it requires computing the singular values with the complexity as: $\mathcal{O}(\left|\mathcal{U}\right|d^{2}+\left|\mathcal{I}\right|d^{2})$.
Proposition 2.
When $f\left(\sigma_{1},\cdots,\sigma_{d}\right)=\sum_{i}\sigma_{i}^{2K}$, $K\in\mathbb{N}_{+}$ and $K\geq 2$, we can balance the spectrum with the following Equation:
(13) $\mathbf{H}\leftarrow\mathbf{H}-\alpha\left(\mathbf{H}\mathbf{H}^{T}\right)^{K-1}\mathbf{H}.$
Input : Embedding matrix $\mathbf{H}$; batch size $B$; hyperparameter $\alpha\in\mathbb{R}^{+}$
sample the users and items from $\mathbf{R}^{+}$: $\mathcal{U}_{B}$, $\mathcal{I}_{B}$ ;
generate the normalized embeddings: $\mathbf{H}_{B}^{U}$, $\mathbf{H}_{B}^{I}$;
$\mathbf{H}_{B}^{U}\leftarrow\mathbf{H}_{B}^{U}-\alpha\mathbf{H}_{B}^{U}{\mathbf{H}_{B}^{U}}^{T}\mathbf{H}_{B}^{U}$;
$\mathbf{H}_{B}^{I}\leftarrow\mathbf{H}_{B}^{I}-\alpha\mathbf{H}_{B}^{I}{\mathbf{H}_{B}^{I}}^{T}\mathbf{H}_{B}^{I}$;
update parameters via $\mathcal{L}(\hat{r}_{ui},r_{ui}),(u,i)\in\mathbf{R}^{+}$
Algorithm 1 DirectSpec
According to SVD: $\mathbf{H}=\mathbf{P}diag(\sigma_{i})\mathbf{Q}^{T}$ with $\mathbf{P}$, $\mathbf{Q}$, and $diag(\cdot)$ as the left, right stacked singular vectors, and the diagonalization operator, respectively. It is obvious that:
(14) $\left(\mathbf{H}\mathbf{H}^{T}\right)^{K}=\left(\mathbf{P}diag\left(\sigma_{i}^{2}\right)\mathbf{P}^{T}\right)^{K}=\mathbf{P}diag\left(\sigma_{i}^{2K}\right)\mathbf{P}^{T},$
due to the orthogonality of singular vectors: $\mathbf{Q}^{T}\mathbf{Q}=\mathbf{I}$, $\mathbf{P}^{T}\mathbf{P}=\mathbf{I}$. Then, we can make the following observation:
(15) $f\left(\sigma_{1},\cdots,\sigma_{d}\right)=\sum_{i}\sigma_{i}^{2K}=\operatorname{Tr}\left(\mathbf{P}diag\left(\sigma_{i}^{2K}\right)\mathbf{P}^{T}\right)=\operatorname{Tr}\left(\left(\mathbf{H}\mathbf{H}^{T}\right)^{K}\right),$
where the trace $\operatorname{Tr}(\cdot)$ is equal to the sum of eigenvalues. Then, the embedding can be updated via SGD as follows:
(16) $\mathbf{H}\leftarrow\mathbf{H}-\alpha\frac{\partial\operatorname{Tr}\left(\left(\mathbf{H}\mathbf{H}^{T}\right)^{K}\right)}{\partial\mathbf{H}}=\mathbf{H}-\alpha K\left(\mathbf{H}\mathbf{H}^{T}\right)^{K-1}\mathbf{H}.$
By simply setting $\alpha\leftarrow\alpha K$, we obtain Equation (13). Note that we can ignore the constraint term (i.e., $\sum_{i}\sigma_{i}^{2}=c^{2}$) as it can be replaced by the normalization: $\mathbf{H}_{i}\leftarrow\frac{\mathbf{H}_{i}}{\left|\mathbf{H}_{i}\right|}$ where $\mathbf{H}_{i}$ represents an arbitrary row of $\mathbf{H}$, simply due to the following relation:
(17) $\sum_{i}\sigma_{i}^{2}=\operatorname{Tr}\left(\mathbf{P}diag\left(\sigma_{i}^{2}\right)\mathbf{P}^{T}\right)=\operatorname{Tr}(\mathbf{H}\mathbf{H}^{T})=\sum_{i}\mathbf{H}_{i}\mathbf{H}_{i}^{T}=d,$
where the diagonal element $\mathbf{H}_{i}\mathbf{H}_{i}^{T}=1$. To further shed light on how Equation (13) balances the spectrum, we can rewrite it as follows:
(18) $\mathcal{F}\left(\mathbf{H}\right)=\mathbf{H}-\alpha\left(\mathbf{H}\mathbf{H}^{T}\right)^{K}\mathbf{H}$ $=\mathbf{P}diag\left(\sigma_{i}\right)\mathbf{Q}^{T}-\alpha\mathbf{P}diag\left(\sigma_{i}^{2K+1}\right)\mathbf{Q}^{T},$
$=\mathbf{P}diag\left(\sigma_{i}\left(1-\alpha\sigma^{2K}_{i}\right)\right)\mathbf{Q}^{T},$
where we redefine $K\geq 1$ for simplicity. As the function $\sigma_{i}^{2K}$ increases faster than $\sigma_{i}$, $1-\alpha\sigma^{2K}_{i}>1-\alpha\sigma^{2K}_{j}$ for $\sigma_{i}<\sigma_{j}$ as long as $\alpha$ is properly set (i.e., $01-\alpha\sigma_{1}^{2K}>$ 0). Thus, $1-\alpha\sigma_{i}^{2K}$ can be considered as rescaled factors: the larger singular values are multiplied by smaller weights, leading to a balanced distribution. Note that each update would not shrink the singular values as the embedding normalization keeps their $L_{2}$ norm unchanged as shown in Equation (17). According to the definition of entropy and erank, a more uniform distribution results in a higher erank, indicating that $erank(\mathcal{F}\left(\mathbf{H}\right))\geq erank(\mathbf{H})$ where $=$ holds when $erank(\mathbf{H})=d$. To further enhance the efficiency of DirectSpec, we set $K=1$ and adopt a batch training strategy instead of updating the whole $\mathbf{H}$. We can integrate DirectSpec as a spectrum-normalization layer in the recommendation algorithm as formulated in Algorithm 1 and as illustrated in Figure 3 (a), where we do not need to explicitly sample the negative pairs, saving the training cost. To demonstrate and validate the effectiveness of DirectSpec, we run a toy example as shown in Figure 3 (b).
Figure 3. (a): An illustration of the proposed DirectSpec, (b): A toy example running DirectSpec on a randomly generated matrix with torch.randn(10,10) $\{5,10\}$ times.
4.2. Spectrum Balancing Could be Poisonous
Proposition 3.
The goal of spectrum balancing and alignment of positive pairs is contradictory.
Suppose the user $u$ and $v$ both interacted with item $i$. Let $\mathbf{H}_{U}$ denote the embedding matrix for users, then $\mathbf{H}_{U}\mathbf{H}_{U}^{T}=\mathbf{I}$ when the perfectly balanced spectrum is achieved: $\sigma_{1}=\cdots=\sigma_{d}$, indicating that $u$ and $v$ are orthogonal in the embedding space: $\mathbf{h}_{u}^{T}\mathbf{h}_{v}=0$. On the other hand, $i$ is optimized to be parallel to both $u$ and $v$: $\mathbf{h}_{u}^{T}\mathbf{h}_{i}=1$ and $\mathbf{h}_{v}^{T}\mathbf{h}_{i}=1$ when the perfect alignment is reached, while such a relation is impossible in Euclidean space: a vector is both parallel to two vectors orthogonal to each other. Therefore, balancing spectrum could be poisonous to alignment of positive pairs and vice versa. In figure 4, we investigate how erank, training loss, and accuracy (nDCG@10) change with the intensity of spectrum balancing by adjusting $\alpha$ of DirectSpec. (i.e., larger (smaller) $\alpha$ indicates intenser (weaker) spectrum balancing). We can see that erank rises as increasing $\alpha$, reaching almost $d$ at $\alpha=0.1$. In the meanwhile, the training loss representing the distance among positive pairs (i.e., lower (higher) loss indicates a better (worse) alignment of positive pairs) accordingly rises with erank in Figure 4 (a), demonstrating the contradiction between spectrum balancing and alignment. The accuracy increases and then drops with $\alpha$, reaching the maximum at $\alpha=0.02$ and erank$=62.9$ in Figure 4 (b), showing that the perfect balanced spectrum does not necessarily lead to better performance. In conclusion, a good representation lies in the trade-off between alignment and spectrum balancing, rather than focusing solely on either of them.
Figure 4. How (a): the effective rank and training loss (b): the effective rank and accuracy change with $\alpha$.
4.3. Individualized Spectrum Balancing (DirectSpec+)
As pointed out in Section 4.2 that balancing the embedding spectrum via DirectSpec is equivalent to the orthogonalization of unobserved pairs with each user/item being optimized as:
(19) $\mathbf{h}_{i}\leftarrow\mathbf{h}_{i}-\sum_{j}\alpha\left(\mathbf{h}_{i}^{T}\mathbf{h}_{j}\right)\mathbf{h}_{j}.$
Compared to $\alpha$, $\mathbf{h}_{i}^{T}\mathbf{h}_{j}$ can be considered as an individualized coefficient controlling the intensity for balancing the spectrum between $i$ and $j$. Given that the goal is not solely to generate a perfectly spectrum but rather to balance the trade-off between alignment and spectrum optimization, we argue that DirectSpec does not fully consider the following two factors:
  • The intensity should be adaptive to the extent of collapse (i.e., the pairs suffering more (less) from the collapse issue should be balanced with a higher (lower) intensity ).
  • Unobserved pairs may also show relations with varying degrees (i.e., the pairs showing closer relations should be balanced with a lower intensity).
We propose an enhanced individualized spectrum balancing design DirectSpec+ as follows:
(20) $\mathbf{H}_{B}^{U}\leftarrow\mathbf{H}_{B}^{U}-\alpha\cdot{\rm softmax}\left(\mathbf{H}_{B}^{U}{\mathbf{H}_{B}^{U}}^{T}\odot\mathbf{\Gamma}\right)\mathbf{H}_{B}^{U},$
$\mathbf{H}_{B}^{I}\leftarrow\mathbf{H}_{B}^{I}-\alpha\cdot{\rm softmax}\left(\mathbf{H}_{B}^{I}{\mathbf{H}_{B}^{I}}^{T}\odot\mathbf{\Gamma}\right)\mathbf{H}_{B}^{I},$
where $\odot$ represents the element-wise multiplication. Here, the individualized coefficient between $i$ and $j$ is replaced by $\frac{\exp(\tau_{ij}\cdot\mathbf{h}_{i}^{T}\mathbf{h}_{j})}{\sum_{k}\exp(\tau_{ik}\cdot\mathbf{h}_{i}^{T}\mathbf{h}_{k})}$, where the pairs suffering more from the collapse (i.e., $\mathbf{h}_{i}^{T}\mathbf{h}_{j}\rightarrow 1$) are balanced with a higher intensity. $\tau_{ij}$ is a temperature reflecting the relations between $i$ and $j$ estimated from the interaction data, and we propose two adaptive temperature designs to adapt to the pairs with varying degrees of correlations:
  • i.
    We dynamically learn the temperature through an attention mechanism such as $\mathbf{\Gamma}_{uv}=\sigma(\mathbf{W}^{T}[\mathbf{h}_{u}||\mathbf{h}_{v}])$ for $u$ and $v$, where ||||| | stands for the concatenate operation and $\mathbf{W}\in\mathbb{R}^{2d}$ is the transform matrix.
  • ii.
    We use a unparameterrized polynomial graph filter $\phi(\cdot)=\sum_{l=0}^{L}\mathbf{\hat{A}}^{l}$ to measure the relation: $\tau_{uv}=\pi(\cos(\phi(u),\phi(i))$), where $\cos(\cdot,\cdot)$ is the cosine similarity. The user/item pairs with a higher similarity indicating a closer relation should be optimized with a smaller temperature (i.e., lower intensity). Thus, we use a linear function $\pi(\cdot)$ mapping the similarity to a proper range $[\tau_{0},\tau_{1}]$: the highest and lowest similarity is mapped to $\tau_{0}$ and $\tau_{1}$, respectively.
We will compare the two designs in Section 5.3.2. We substitute Equation (20) into Algorithm 1 and obtain the running algorithm for DirectSpec+.
4.4. Discussion
4.4.1. Connection with SSL-based Methods
Self-supervised learning (SSL) has shown great potential in recent years. Some SSL methods can generate high-quality representations without explicitly sampling negative samples that have been applied to recommendation as well, such as Barlow Twins (BT) (Zbontar et al., 2021), LogDet (Chen et al., 2023), and spectral contrastive loss (SCL) (HaoChen et al., 2021):
(21) $\mathcal{L}_{BT}=\sum_{i}\left(1-\mathcal{C}_{ii}\right)^{2}+\lambda\sum_{i}\sum_{j\neq i}\mathcal{C}_{ij}^{2},$
$\mathcal{L}_{logdet}=\operatorname{Tr}\left(\Sigma\right)-\log\det\left(\Sigma\right),$
$\mathcal{L}_{SCL}=\sum_{i}\sum_{i\neq j}\left(\mathbf{h}_{i}^{T}\mathbf{h}_{j}\right)^{2},$
where $\mathcal{C}_{ij}=\mathbf{H}[:,i]^{T}\mathbf{H}[:,j]$ is the covariance, and $\Sigma=\mathbf{H}^{T}\mathbf{H}$ is the covariance matrix.
Proposition 4.
BT, LogDet, and SCL are special cases of DirectSpec (Equation (11)) with $f\left(\sigma_{1},\cdots,\sigma_{d}\right)=\sum_{i}(\sigma_{i}^{2}-1)^{2}$ for BT and SCL and $f\left(\sigma_{1},\cdots,\sigma_{d}\right)=\sum_{i}\sigma_{i}-2\log\sigma_{i}$ for LogDet, respectively.
With $\mathbf{H}$ being normalized column-wisely, $\sum\sigma_{i}^{2}=\operatorname{Tr}(\mathbf{Q}diag(\sigma_{i}^{2})\mathbf{Q}^{T})=\operatorname{Tr}(\mathbf{H}^{T}\mathbf{H})=d$. Then, Barlow Twins can be rewritten as follows:
(22) $\mathcal{L}_{BT}$ $=\operatorname{Tr}\left(\left(\mathbf{H}^{T}\mathbf{H}-\mathbf{I}\right)^{2}\right)=\operatorname{Tr}\left(\mathbf{Q}diag\left(\sigma_{i}^{2}\right)\mathbf{Q}^{T}-2\mathbf{Q}diag\left(\sigma_{i}\right)\mathbf{Q}^{T}+\mathbf{Q}\mathbf{Q}^{T}\right)$
$=\operatorname{Tr}\left(\mathbf{Q}diag\left((\sigma_{i}^{2}-1)^{2}\right)\mathbf{Q}^{T}\right)=\sum_{i}\left(\sigma_{i}^{2}-1\right)^{2}.$
With $\mathbf{H}$ being normalized row-wisely, we can rewrite SCL as a similar form to BT:
(23) $\mathcal{L}_{SCL}=\operatorname{Tr}\left(\left(\mathbf{H}\mathbf{H}^{T}-\mathbf{I}\right)^{2}\right)=\operatorname{Tr}\left(\mathbf{P}diag\left((\sigma_{i}^{2}-1)^{2}\right)\mathbf{P}^{T}\right)=\sum_{i}\left(\sigma_{i}^{2}-1\right)^{2}.$
Similarly, we can reformulate LogDet as:
(24) $\mathcal{L}_{logdet}=\operatorname{Tr}\left(\mathbf{H}^{T}\mathbf{H}\right)-\log\det\left(\mathbf{H}^{T}\mathbf{H}\right)=\sum_{i}\sigma_{i}^{2}-\log\prod\sigma_{i}^{2}=\sum_{i}\sigma_{i}^{2}-2\log\sigma_{i}.$
By substitute Equation (22) (or (23)) and (24) into Equation (11) and solve them, we can get the minima achieved at $\sigma_{1}=\cdots=\sigma_{d}=\sqrt{d}$ for the three methods, $\beta=1-2d$ and $\beta=\frac{1-d}{d}$ for BT/SCL and LogDet when $c=d$, respectively, demonstrating that they can prevent the representations from collapsing by implicitly balancing the embedding spectrum. Since they belong to the DirectSpec framework, our proposed DirectSpec+ balances spectrum more adaptively and effectively, showing superiority over them. It is worth noting that InfoNCE loss cannot be reformulated as a case of DirectSpec. However, the embeddings are nearly and tend to be more orthogonal to each other as the dimensionality $d$ becomes larger, indicating a perfectly balanced spectrum, as shown in Proposition 3 of (Wang and Isola, 2020). We will compare our proposed methods with them in Section 5. Given DirectSpec’s generality and similarity to SSL methods that have been applied across various domains, we believe that DirectSpec can also contribute to other domains such as images, graph-structured data, etc.
Table 1. Comparison of time complexity.
Model DirectSpec Barlow Twins SCL LogDet
Complexity $c|\mathbf{R}^{+}|Bd$ $c|\mathbf{R}^{+}|d^{2}$ $c|\mathbf{R}^{+}|Bd$ $c|\mathbf{R}^{+}|d^{2}\left(1+\frac{d}{B}\right)$
Table 2. Statistics of datasets.
Datasets #User #Item #Interactions Density%
CiteULike 5,551 16,981 210,537 0.223
Yelp 25,677 25,815 731,672 0.109
Gowalla 29,858 40,981 1,027,370 0.084
4.4.2. Complexity
We compare the complexity of DirectSpec+ with Barlow Twins, spectral contrastive loss, and LetDet in Table 1, where $B$ and $c$ denote the batch size and the required epochs, respectively. We ignore the complexity for optimizing the alignment here. Overall, LogDet has the most expensive computational cost as it requires computing the matrix determinant with complexity as $\mathcal{O}(d^{3})$. The complexity of DirectSpec+ is identical to that of SCL. The difference of complexity between DirectSpec+ and BT lies in the term $Bd$ and $d^{2}$, where DirectSpec+ might be more costly when $B$ is large. However, considering that we set $B=256$ in the experiments, and the complexity of back propagation for BT and SCL is still $c|\mathbf{R}^{+}|Bd$ which is not contained in Table 1, DirectSpec+ is more time-efficient than BT and SCL.
4.4.3. Comparison between Embedding Collapse and Over-Smoothing
Generally, embedding collapse can be mathematically formulated as follows:
(25) ${\rm rank}\left(\mathbf{H}^{(L)}\right)<\epsilon,$
where $\mathbf{H}^{(L)}$ is the embedding when erank converges after updating $L$ times, $\epsilon<d$, and a smaller $\epsilon$ indicates more serious embedding collapse. Over-smoothing is attributed to the shrunk graph spectrum as repeating the message passing that can also be formulated as Equation (25). The collapse issue analyzed in this work shares similarities with over-smoothing, due to the observation that the parameter-updating process in recommendation algorithm shares the same form as the message passing in GNN. However, since the cause of over-smoothing in GNN lies in the graph spectrum, existing works tackle this issue mostly by dealing with the graph spectrum, such as randomly dropping out the edges to perturb the spectrum (Rong et al., 2020), balancing the contribution of different order neighborhood (Gasteiger et al., 2019; Peng et al., 2022a) to avoid all eigenvalues from shrinking, etc., which fail to solve the embedding collapse in this work, as it exists generally on recommendation algorithms without relying on explicit graph structures. On the other hand, our proposed DirectSpec tackles a general embedding collapse by directly balancing the embedding spectrum which is proved to solve this issue. We will show that our DirectSpec is also effective on over-smoothing and embedding collapse caused by other reasons in Section 5.
4.4.4. Comparison with Related Works
Recently, several efforts have been made to address ‘collapse issues’ in recommendation methods (Chen et al., 2023; Zhang et al., 2023; Guo et al., 2024), which shares similarity with our work. Guo et al. (Guo et al., 2023) points out an embedding collapse within the feature interaction module for click-rate prediction (CTR) caused by the imbalanced gradients over different singular vectors, while the embedding collapse issue in our work stems from optimizing a CF algorithm. The dimensional collapse in graph collaborative filtering (GCF) investigated in (Chen et al., 2023; Zhang et al., 2023) overlap more significantly with ours. Both the works exploit a LogDet function to address the issue, which are more computationally expensive than DirectSpec as analyzed in Section 4.4.2. We will demonstrate the superiority of our proposed methods in Section 5.
5. Experiments
In this section, we comprehensively evaluate our proposed methods on three public datasets. We implement DirectSpec and DirectSpec+ on two popular baselines: MF and LightGCN. Particularly, we introduce data description and implementation details in Section 5.1; we compare the proposed methods with other competitive baselines in terms of accuracy and efficiency in Section 5.2. Finally, we conduct model analysis in Section 5.3, including detailed analysis and experimental results on how DirectSpec+ prevents embeddings from collapsing without relying on negative samples, and how different settings of hyperparameters affect model performance.
5.1. Experimental Settings
5.1.1. Datasets
We use the following three publicly available real-world datasets in this work, where the statistics of them are summarized in Table LABEL:datasets.
  • CiteULike [^1] https://github.com/js05212/citeulike-a: This is an scholarly article recommendation dataset. Users are allowed to create their own collections of articles including abstracts, titles, and tags, etc.
  • Yelp (He et al., 2016): This is a business dataset from Yelp Challenge data. The items are point of interests (POIs), users can leave reviews and ratings.
  • Gowalla (Wang et al., 2019): This is a check-in datasets recording which locations users have visited.
Since we focus on CF for implicit feedbacks, we remove auxiliary information such as reviews, tags, geological and item information, etc, and transform explicit ratings to 0-1 implicit feedbacks.
5.1.2. Evaluation Metrics
We adopt two widely used evaluation metrics for personalized rankings: Recall and normalized discounted cumulative gain (nDCG) (Järvelin and Kekäläinen, 2002) to evaluate model performance. The recommendation list is generated by ranking unobserved items and truncating at position $k$. Recall measures the ratio of the relevant items in the recommendation list to all relevant items in test sets, while nDCG takes the rank into consideration by assigning higher scores to the relevant items ranked higher. We use 80% of the interactions for training and randomly select 10% from the training set as validation set for hyper-parameter tuning, the rest 20% data is used for test sets; we report the average accuracy on test sets.
5.1.3. Baselines
We compare DirectSpec with the following competitive baselines:
  • BPR (Rendle et al., 2009): This is a stable and classic MF-based method, exploiting a Bayesian personalized ranking loss for personalized rankings.
  • ENMF (Chen et al., 2020b): This is a neural recommendation model performing whole-data-based learning without sampling. We choose user-based ENMF as the baseline.
  • CCL (Mao et al., 2021): The proposed cosine contrastive loss (CCL) maximizes the similarity between positive pairs and minimizes the similarity of negative pairs below the margin $m$. After parameter tuning, we set $\mathcal{N}=1000$, $w=300$ on all datasets, $m=0.1$, $0.3$, and $0.3$ on CiteULike, Yelp, and Gowalla, respectively, and we choose MF as the encoder.
  • DirectAU (Wang et al., 2022): This method directly optimizes alignment and uniformity and shows superior performance. Following the original paper, we choose MF and LightGCN as the encoder, and set $\gamma=5.0$, 0.5, and 1.5 on CiteULike, Yelp, and Gowalla, respectively.
  • LightGCN (He et al., 2020): This is a linear GNN method that only keeps neighborhood aggregation for recommendation. We employ a three-layer architecture as our baseline.
  • SGL-ED (Wu et al., 2021): This model explores self-supervised learning by maximizing the agreements of multiple views from the same node, where the node views are generated by adding noise such as randomly removing the edges or nodes on the original graph. We set $\tau=0.2$, $\lambda_{1}=0.1$, $p=0.1$, and use a three-layer architecture.
  • LightGCL (Cai et al., 2023): This is a simple yet effective graph contrastive learning method injecting global collaborative relations via singular value decomposition.
  • GDE (Peng et al., 2022a): This method only keeps a very few graph features for recommendation without stacking layers, showing less complexity and higher efficiency than GNN-based methods.
  • LogDet (Zhang et al., 2023): This is a decorrelation-enhanced method to mitigate the dimensional collapse of graph collaborative filtering.
  • AFDGCF (Wu et al., 2024): This method tackles the over-correlation issue in GNN-based methods, and shares significant similarities with Barlow Twin (Zbontar et al., 2021). It can be considered as the implementation of Barlow Twin for recommendation. We choose LightGCN as the encoder.
5.1.4. Implementation Details
We implemented our DirectSpec based on PyTorch, and the code is released publicly [^2] https://github.com/tanatosuu/directspec. SGD is adopted as the optimizer for all models, the embedding size $d$ is set to 64, the regularization rate is set to 0.01 on all datasets, the learning rate is tuned amongst $\{0.001,0.005,0.01,\cdots\}$; without specification, the model parameters are initialized with Xavier Initialization (Glorot and Bengio, 2010); the batch size is set to 256. We report other hyperparameter settings in the next subsection.
Table 3. Performance comparison. The best baseline is underlined. “*” indicates statistical significance at $p<0.01$ for a one-tailed t-test.
Datasets Methods nDCG@10 nDCG@20 Recall@10 Recall@20
CiteULike BPR 0.1620 0.1773 0.1778 0.2190
CCL(MF) 0.1545 0.1642 0.1716 0.1996
ENMF-U 0.1335 0.1601 0.1556 0.2193
DirectAU(MF) 0.2102 0.2252 0.2260 0.2693
DirectAU(GCN) 0.1926 0.2109 0.2116 0.2604
LightGCN 0.1610 0.1781 0.1771 0.2190
SGL-ED 0.1890 0.2065 0.2117 0.2588
LightGCL 0.2096 0.2238 0.2214 0.2638
GDE 0.1890 0.2061 0.2055 0.2528
LogDet 0.1403 0.1579 0.1548 0.2034
AFDGCF 0.1619 0.1787 0.1790 0.2246
DirectSpec(MF) 0.1688 0.1849 0.1827 0.2270
DirectSpec(GCN) 0.1693 0.1863 0.1875 0.2334
DirectSpec+(MF) 0.2197 0.2354 0.2352 0.2818
DirectSpec+(GCN) 0.2038 0.2213 0.2213 0.2704
Yelp BPR 0.0487 0.0583 0.0607 0.0869
CCF(MF) 0.0509 0.0593 0.0617 0.0846
ENMF-U 0.0665 0.0811 0.0786 0.1167
DirectAU(MF) 0.0721 0.0848 0.0872 0.1219
DirectAU(GCN) 0.0695 0.0819 0.0854 0.1194
LightGCN 0.0572 0.0690 0.0721 0.1045
SGL-ED 0.0676 0.0794 0.0837 0.1166
LightGCL 0.0673 0.0790 0.0836 0.1151
GDE 0.0653 0.0771 0.0805 0.1129
LogDet 0.0711 0.0833 0.0863 0.1203
AFDGCF 0.0585 0.0702 0.0743 0.1065
DirectSpec(MF) 0.0689 0.0804 0.0839 0.1156
DirectSpec(GCN) 0.0712 0.0832 0.0861 0.1192
DirectSpec+(MF) 0.0743 0.0864 0.0909 0.1249
DirectSpec+(GCN) 0.0745 0.0868 0.0909 0.1252
Gowalla BPR 0.1164 0.1255 0.1186 0.1483
CCL(MF) 0.1269 0.1349 0.1295 0.1573
ENMF-U 0.1166 0.1286 0.1173 0.1528
DirectAU(MF) 0.1286 0.1402 0.1349 0.1710
DirectAU(GCN) 0.1298 0.1409 0.1363 0.1711
LightGCN 0.0987 0.1098 0.1074 0.1399
SGL-ED 0.1343 0.1462 0.1417 0.1779
LightGCL 0.1368 0.1482 0.1436 0.1801
GDE 0.1261 0.1367 0.1313 0.1656
LogDet 0.1179 0.1272 0.1203 0.1511
AFDGCF 0.1011 0.1117 0.1087 0.1406
DirectSpec(MF) 0.1133 0.1225 0.1183 0.1480
DirectSpec(GCN) 0.1160 0.1251 0.1191 0.1490
DirectSpec+(MF) 0.1389 0.1509 0.1447 0.1819
DirectSpec+(GCN) 0.1405 0.1518 0.1467 0.1837
Table 4. Training time (seconds) per epoch.
Model/Data Citeulike Yelp Gowalla
BPR 2.37 8.53 13.28
DirectSpec+(MF) 3.78 (1.59x) 13.15 (1.54x) 19.10 (1.44x)
LigtGCN 8.78 43.18 79.79
DirectSpec+(GCN) 9.49 (1.08x) 45.70 (1.06x) 81.20 (1.02x)
SGL-ED(GCN) 22.20 (2.53x) 182.54 (4.23x) 401.96 (5.04x)
LightGCL(GCN) 11.59 (1.32x) 51.38 (1.19x) 92.56 (1.16x)
AFDGCF(GCN) 11.45 (1.30x) 48.53 (1.12x) 84.86 (1.06x)
5.2. Comparison
5.2.1. Performance
We report overall performance in Table 3, and observe the followings:
  • Overall, GNN-based methods show better performance on sparse data (Yelp and Gowalla) than dense data (CiteULike). For instance, DirectSpec+(GCN) achieves better and worse than DirectSpec+(MF) on Gowalla and CiteULike, respectively, and their performance is close on Yelp. LightGCN underperforms BPR on CiteULike and Gowalla, which might be attributed to the slow convergence that has been reported in the original paper. Among baselines, DirectAU achieves the best on CiteULike and Yelp, while LightGCL outperforms other baselines on Gowalla. Our proposed DirectSpec+ implemented on both LightGCN and MF consistently show improvements over all baselines.
  • ENMF and CCL exploit enhanced negative sampling for training, while the superior performance over them demonstrates the superiority of DirectSpec learning paradigm.
  • Unlike negative sampling, SSL-based methods adopt different training strategies to enhance the data representations. Amongst them, DirectAU directly regulating the uniformity shows relatively superior performance, while AGDGCF utilizing a Barlow-Twin-like model to remove the feature correlation shows poor performance.
  • Since our methods are implemented on MF and LightGCN, the effectiveness of our methods can be further demonstrated by directly comparing with them. DirectSpec(MF) and DirectSpec+(MF) outperform BPR by 14.3% and 35.8%, on average in terms of nDCG@10, respectively. In the meanwhile, the improvement of DirectSpec(GCN) and DirectSpec+(GCN) over LightGCN is 15.7% and 33.1% on average in terms of nDCG@10, respectively.
  • Overall, the improvement of DirectSpec over MF and LightGCN is Yelp$>$Gowalla$>$CiteULike (33.0%, 7.4%, and 4.7% on average in terms of nDCG@10, respectively). Figure 5 compares the extent of collapse on three datasets using MF and LightGCN (with only positive samples). We can observe that Yelp suffers more from the collapse while CiteULike suffers less, which is consistent with the improvement of DirectSpec. This observation reveals that the dataset suffering more from collapse tends to benefit more from DirectSpec.
  • DirectSpec+ shows significant improvement over DirectSpec across all datasets. As introduced in Section 4.3, we adopt a individualized spectrum balancing policy for DirectSpec+ to achieve a better trade-off between alignment and spectrum balancing. The superior performance demonstrates its effectiveness.
Figure 5. The extent of collapse on three datasets.
(a) CiteULike
(b) Yelp
(c) CiteULike
(d) Yelp
Figure 6. All models are optimized without using observed interactions. In (a) and (b), we show how erank changes on DirectSpec (MF) and MF (BCE) where $N$ is the negative sampling ratio. In (c) and (d), we use BCE (MF) for the first 50 epochs and DirectSpec for the last 50 epochs.
5.2.2. Efficiency
We report the training time per epoch of several methods implemented on BPR and LightGCN in Table 4. The experiments are all conducted on Intel(R) Core(TM) i9-10980XE CPU and NVIDIA RTX A6000 GPU. We first compare BPR and DirectSpec+(MF), and observe that it takes roughly 0.5x additional running time, which is acceptable considering the significant improvement. Compared to other methods implemented on LightGCN, DirectSpec+ almost shows no additional running time, due to the reason that the complexity of DirectSpec+ is mainly determined by the batch size rather than the size of users and items. Since LightGCN has a higher complexity than BPR, the proportion of DirectSpec+(LightGCN)’s running time per epoch to that of LightGCN is much smaller than the proportion of DirectSpec+(MF)’s running time to that of BPR. This finding can explain another observation that the additional time is smaller on larger datasets (i.e., Gowalla$<$Yelp$<$CiteULike). The complexity of DirectSpec+ is unchanged while LightGCN and BPR accordingly take more running time on larger datasets, thus the proportion of DirectSpec+’s running time to that of LightGCN/BPR becomes smaller. Meanwhile, SGL-ED is more computationally expensive on larger datasets, and DirectSpec+(LightGCN) is more efficient than it with more significant improvement over LightGCN. Furthermore, despite the additional training time per epoch added to the encoder, the scalability and utility have actually been improved. For instance, the required epochs are 250 and 600 for BPR and LightGCN on Gowalla, respectively, as opposed to just 80 and 60 epochs for DirectSpec(MF) and DirectSpec(LightGCN), resulting in 2.2x and 9.8x speed-up, respectively.
(a) AFDGCF on CiteULike
(b) LogDet on Gowalla
(c) DropEdge on Yelp
Figure 7. Comparison with several methods.
(a) Erank comparison
(b) Singular value distribution of the embedding matrix
(c) Comparison of erank and average similarity on pairs in test sets
Figure 8. How DirectSpec prevents collapse (on Yelp).
5.2.3. Comparison with Negative Sampling
The results in Table 3 demonstrates the superiority of DirectSpec over negative sampling based methods including BPR and CCL in terms of performance. Furthermore, we compare how these two types of methods alleviate collapse issue by optimizing them solely based on negative pairs. In Figure 6 (a) and (b), we observe that the erank of MF (BCE) consistently drops without optimization on the positive pairs, resulting in a complete collapse, and the erank drops more quickly as increasing the negative sampling ratios, whereas the erank of DirectSpec remains unchanged throughout the training. This observation indicates that there exists a dynamic balance between positive and negative samplings, the lack of either of them cannot prevent the embeddings from collapsing. Furthermore, since negative sampling is equivalent to a high pass filter, it can only alleviate the collapse caused by low pass filters (e.g., optimization on observed interactions), while DirectSpec is equivalent to an all pass filter that can tackle the collapse under any situations. Figure 6 (c) and (d) demonstrate that DirectSpec can recover the embeddings from collapsing caused by negative sampling.
5.2.4. Comparison with SSL-based methods
Particularly, we focus on AFDGCF (BT) and LogDet that can be considered as special cases of DirectSpec. We show how accuracy and erank change with the hyperparameters $\alpha$ and $\lambda$ controlling the intensity of spectrum balancing in Figure 7 (a) and (b), respectively. We observe that both performance and erank are not sensitive to $\alpha$ on AFDGCF, which are significantly inferior to DirectSpec+. While the erank consistently increases with $\lambda$ and reaches almost $d$ at $\lambda=0.9$ on LogDet, the accuracy drops early at $\lambda=0.3$ with erank=61.64. On the other hand, the best performance of DirectSpec+ is achieved at erank=63.77, showing significant improvement over LogDet. This observation indicates that DirectSpec+ can balance the spectrum in a more reasonable way, achieving a better trade-off between alignment and spectrum balancing than LogDet and AFDGCF.
5.2.5. Comparison with Solutions to Overs-Smoothing
[…]
(a) CiteULike
(b) Yelp
Figure 9. Impact of $\alpha$ on accuracy (nDCG@10) and erank.
(a) CiteULike
(b) Yelp
(c) Gwoalla
(d) Gowalla
Figure 10. (a) and (b) show how the accuracy changes with $\tau_{1}$, (c) displays the sensitivity of $K$, and (d) illustrates how the accuracy changes with $\tau_{0}$.
5.3. Model Analysis
5.3.1. Can DirectSpec Prevent Collapse?
In Figure 8 (a), we report how the erank changes as training proceeds on BPR, LightGCN, and DirectSpec+(MF). For BPR and LightGCN, the erank plunges significantly in the first place, then increases slowly and tends to converge to a value smaller than $d$, resulting in an incomplete embedding collapse. While the erank of DirectSpec+(MF) is close to $d$ at the beginning of training and remains almost unchanged throughout the training. Figure 8 (b) illustrates the normalized singular value distribution of the embedding matrix when the accuracy of BPR and DirectSpec+(MF) is maximized. We can see that most singular values of BPR’s embeddings are less than 0.2, barely contributing to the representations; while DirectSpec+(MF)’s singular values are mostly larger than 0.5, indicating that the representations are contributed by more dimensions.
5.3.2. DirectSpec vs. DirectSpec+
The goal of DirectSpec+ is not simply to generate a perfectly balanced spectrum like DirectSpec, but to balance the spectrum while also preserving the mapped relationships among users/items in the embedding space. In Figure 8 (c), the average similarity on pairs in test sets are defined as: $SIM=\frac{\sum_{u}\frac{\sum_{i\in\mathcal{N}^{t}_{u}}\sigma(\mathbf{h}_{u}^{T}\mathbf{h}_{i})}{|\mathcal{N}^{t}_{u}|}}{|\mathcal{U}|}$ where $\mathcal{N}^{t}_{u}$ is the item set $u$ has interacted with in test sets and $\sigma(\cdot)$ is the sigmoid function. In spite of a higher erank DirectSpec has shown during training, DirectSpec+ still outperforms DirectSpec as it shows a higher average similarity on test sets, indicating that DirectSpec+ is better at representing the user/item relationships in the embedding space in a reasonable way.
Table 5. Comparison between static and dynamic temperature design.
Citeulike Yelp Gowalla
nDCG@10 Recall@10 nDCG@10 Recall@10 nDCG@10 Recall@10
Static (i) 0.2197 0.2352 0.0743 0.0909 0.1389 0.1447
Dynamic (ii) 0.1593 0.1693 0.0485 0.0611 0.1126 0.1144
5.3.3. Impact of $\alpha$
$\alpha$ serves as a global coefficient controlling the extent of spectrum balancing. A larger (smaller) $\alpha$ implies that large singular values are multiplied by smaller (larger) weights, leading to a more (less) uniform spectrum distribution. According to Equation (18), DirectSpec could fail when $\alpha$ is set too large: $01-\alpha\sigma_{i}^{2}<$ 0 (with $K=1$). Since it takes additional time to calculate $\sigma_{1}$ during each training, we use the grid search for optimal $\alpha$. As shown in Figure 9, both the accuracy and erank first increase and then drop, showing a similar trend. Specifically, the speed of collapse tends to outrun the speed of spectrum balancing when $\alpha$ is small as erank converges to a value much smaller than $d$. While when $\alpha$ is large, the large singular values in the original distribution are multiplied by too small weights, which might cause a sharp singular value distribution similar to the original one, explaining why erank declines as consistently increasing $\alpha$. In addition, the best performance is achieved at $\alpha=1.1$ on CiteULike and $\alpha=0.6$ on Yelp, and the erank and accuracy are more sensitive when $\alpha$ is relatively small on Yelp, which might be related to the fact that Yelp suffers more from the collapse issue than other datasets.
5.3.4. Adaptive Temperature Design
We first compare the dynamic and static designs proposed in Section 4.3, and report the results in Table 5. We observe that the static design outperforms the dynamic design by a large margin on all datasets, and the reason might be attributed to the data scarcity on CF task, making the algorithm difficult to learn the inherent user/item relations from the interactions data in a dynamic way. Therefore, DirectSpec+ is implemented with a static temperature design, which avoids bringing additional complexity.
For the static temperature design, we first fixed $\tau_{0}=\tau_{1}$ and only tune $\tau_{1}$, then we gradually reduce $\tau_{0}<\tau_{1}$ and observe how accuracy changes accordingly. As shown in Figure 10 (a) and (b), the accuracy first increases and then drops as $\tau_{1}$ increases, which is maximized at $\tau_{1}=3$ on CiteULike and $\tau_{1}=4$ on Yelp. $\tau_{1}$ controls the strength of penalties on the pairs severely suffering from the collapse issue (i.e., highly correlated), the samples that are not highly correlated tend to be ignored when $\tau_{1}$ is set too large, while a too small $\tau_{1}$ fails to optimize the highly correlated pairs in an effective way (DirectSpec+ degenerates to DirectSpec when $\tau_{1}\rightarrow 0$). We take the relations among unobserved pairs into consideration via a polynomial graph filter $\sum_{l=0}^{L}\mathbf{\hat{A}}^{l}$ as reducing $\tau_{0}<\tau_{1}$ (i.e., the pairs showing closer relations are optimized with a lower intensity$<\tau_{1}$). In Figure 10 (c), DirectSpec+ achieves the best performance at $L=3$, incorporating neighborhood farther than $L>3$ results in degradation as they are too far away on the graph. In Figure 10 (d), the accuracy increases and then drops as reducing $\tau$, the maximum is reached at $\tau=3.0$ on Gowalla, demonstrating it effectiveness. We also observe that the improvement from adjusting $\tau_{0}$ is less significant on the other two datasets, which might be related to the data density, as users/items are less hierarchically distributed on denser data. Therefore, we can simple fix: $\tau_{0}=\tau_{1}$ for such datasets.
5.3.5. What Factors Affect DirectSpec’s Performance
As analyzed previously, DirectSpec’s performance is primarily influenced by the extent of collapse and the choice of encoders (i.e., the algorithm implemented on). The variation in improvement caused by different degrees of collapse is unavoidable as our goal is to eliminate the collapse issue. We can choose the encoder according to the data density: GNN-based methods perform better on sparse data while MF-based methods excel in dense data.
6. Related Work
6.1. Collaborative Filtering
Collaborative Filtering (CF), a fundamental task for recommender systems, makes predictions based on user’s past behaviours. Early CF methods exploit users that share similar interests or items that tend to be interacted by similar users to infer user preference (Sarwar et al., 2001). Nowadays, model-based methods are prevalent in CF and recommender systems. The core idea is to map the high dimensional sparse recommendation data to a low-dimensional space, where users and items are characterized as learnable vectors and are optimized based on observed interactions (Koren et al., 2009). With the development of computer hardware and communication network, more information can be collected to analyze user behaviour with advanced algorithms. Besides user-item interactions, side information such as social relations (Jiang et al., 2012), reviews (Bao et al., 2014), geological information (Lian et al., 2014; Peng et al., 2018), and so on can help the algorithms better understand user taste and preference. On another line, early simple CF methods (Koren et al., 2009; Ning and Karypis, 2011) are replaced by more powerful algorithms such as neural network (Sedhain et al., 2015; He et al., 2017), attention mechanism (Chen et al., 2017; Kang and McAuley, 2018), transformer (Sun et al., 2019), etc. Recently, graph neural network (GNN) has also shown tremendous success in recommender systems (Wang et al., 2019; Peng et al., 2022b, 2024). By representing interactions as a bipartite graph, the core idea of GNN is to exploit higher-order neighbor connections to facilitate user/item representations (Kipf and Welling, 2017). However, it has been reported that multi-layer GNNs suffer from an over-smoothing issue that causes user/item representations to be indistinguishable (Peng et al., 2022a; Li et al., 2018). We showed that an embedding collapse issue resembling the over-smoothing exists on recommendation algorithms in general, and our proposed approach has also been shown effective tackling over-smoothing.
6.2. Collapse in Representation Learning
Representation learning has attracted tremendous attention in various research fields (Mikolov et al., 2013; Hjelm et al., 2019), the goal of which is to represent data in an effective and efficient way to make it easier to extract useful information when building predictors. However, it has also been shown that the model tends to collapse where all inputs are mapped to the same constant vector, when only optimizing the model based on the positive pairs (Chen and He, 2021). This issue can be well alleviated by self-supervised learning and contrastive learning by exploiting negative samples in an effective way (Hjelm et al., 2019; Chen et al., 2020a). Due to the heavy computation cost, some research effort has been made to simplify the algorithms without explicitly sampling negative data (Chen and He, 2021; Grill et al., 2020; Zbontar et al., 2021; Zhang et al., 2023). Unfortunately, subsequent works show that a dimensional collapse cannot be ignored where the embedding vectors end up spanning a lower dimensional subspace of the whole embedding space (Hua et al., 2021; Jing et al., 2022). The collapse issue in representation learning shares similarities to some issues such as over-smoothing in GNN (Li et al., 2018), which inspires some researchers to tackle issues in GNN (Guo et al., 2023).
[…]
7. Conclusion and Limitation
In this work, we showed that existing CF methods mostly suffer from an embedding collapse issue. Particularly, alignment of positive pairs is equivalent to a low pass filter, causing the representations of users and items to collapse to a constant vector. Exploiting negative pairs (i.e., negative sampling) can decelerate collapse by acting as a high pass filter to balance the embedding spectrum, while the effectiveness is only limited to certain loss functions which still results in an incomplete collapse, where embeddings are distributed along certain dimensions instead of making full use of all dimensions. To tackle this issue, we proposed a DirectSpec which directly optimizes the spectrum distribution to ensure that all dimensions can contribute to the user/item embeddings as equally as possible. We further pointed out that the key of high-quality representation lies in trade-off between alignment and spectrum balancing, and proposed an enhanced algorithm DirectSpec+, which can balance the spectrum more adaptively and individually. Moreover, we established the connection with SSL-based methods without explicit negative sampling such as Barlow-Twin, LogDet, and spectral contrastive loss by showing that they can be considered as special cases of DirectSpec. Finally, extensive experiments on three publicly available datasets demonstrated the effectiveness and efficiency of our proposed methods. Given the generality of our DirectSpec learning paradigm, we believe that our proposed method can effectively address collapse in representation learning more broadly.
We acknowledge that our proposed methods are still far from perfect, despite the efficiency and effectiveness. Firstly, balancing the embedding spectrum requires tuning hyperparameters such as $\alpha$ and $\tau$, reducing the utility. Furthermore, the effectiveness of some designs (e.g., $\tau_{0}$) is not significant and varies on datasets. We aim to address above limitations in the future work.
Acknowledgement
This paper is based on results obtained from the project, “Research and Development Project of the Enhanced Infrastructures for Post-5G Information and Communication Systems” (JPNP20017), commissioned by the New Energy and Industrial Technology Development Organization (NEDO).